OpenStudy (grant330sims):

Part 1: Explain, in complete sentences, which method you would use to solve the following system of equations. (1 point) Part 2: Explain why you chose that method (1 point) Part 3: Provide the solution to the system. (2 points) x – 3y + 2z = –12 x + 2y + 3z = 6 2x – 3y – z = –2

OpenStudy (anonymous):

What math are you taking?

OpenStudy (grant330sims):

Algebra 2

OpenStudy (anonymous):

Yeah i don't know. I would say substitution.

OpenStudy (grant330sims):

well, I've got how to with two equations, just cant remember how to do three.

OpenStudy (grant330sims):

Having trouble, cuz if you can just tell me what do do, you dont have to give me the answer, I'll just figure it out

OpenStudy (anonymous):

such question is a system of linear equation for further mathematics....so we can reduce it to triangular system and use back substitution...........gaussian elimination would also be appropriate.....or use matrix format with row reduction appplication...actually i am doing an assignment on it

OpenStudy (anonymous):

the y is because we have 3 equations with 3 unknowns in it.so the given methods will be right

OpenStudy (grant330sims):

Bro, i have no idea what you just said..

OpenStudy (anonymous):

no bro but sis......i`ll give u the steps....you follow then do it....ok

OpenStudy (aroub):

Yeah I would say substitution too. Solving the first equation for x in terms of y and z yields, x=3y-2z-12. Substituting the value of x in the second and third equations yields a system of two equations in two unknowns: (3y-2z-12)+2y+3z=6 2(3y-2z-12)-3y-z=-2 => 3y-2z-12+2y+3z=6 6y-4z-24-3y-z=-2 => 5y+z=18 3y-5z=22 for z gives, z=-5y+18 substituting the value of z in the second equation gives: 3y-5(-5y+18)=22 3y+25y-90=22 28y=112 y=4 substituting the value of y in the first equation: 5(4)+z=18 20+z=18 z=-2 substituting the value of y and z in the value of x that we first got ^^ : x=3y-2z-12 x=3(4)-2(-2)-12 x=12+4-12 x=4 The solution set is: { (4,-2,4) }

OpenStudy (grant330sims):

Thank you soo much!

OpenStudy (aroub):

You're welcome!